Exercise 11-08 a-b Suppose the following financial data were reported by 3M Company for 2019 and 2020 (dollars in millions) 3Aint nobody got time for that!
Answer #1

1. Current ratio = current assets / current liability

2019 = 9616 / 5893 = 1.63

2020 = 11161 / 4853 = 2.30

Working Capital = current assets - current liability

2019 = 9616 - 5893 = $3723

2020 = 11161 - 4853 = $6308

2.

Current ratio = current assets / current liability

2020 = (11161 - 197) / (4853 -197) = 2.35

Working Capital = current assets - current liability

2020 = (11161 - 197) - (4853 -197) = $6308